OpenStudy (anonymous):

6x+3y=15 .What does y equal if x=2 ?if x=5

OpenStudy (anonymous):

Just plug back in to the eq'n. y =1 if x=2 and y=-5 if x=5.

OpenStudy (anonymous):

To get there, start by solving the equation for y. 6x+3y=15 3y=15-6x y=5-2x. Then plugging in values of x is much easier.